OpenStudy (anonymous):

A river flows at 1.5 kmph and a boatman who can row at 2.5 kmph on still water, takes 7hours to go a certain distance up the stream and return again to the starting point. Find the total distance covered by the boatman.

OpenStudy (cwrw238):

Hint : going up the stream his speed will be 2.5 - 1.5 = 1 kmph as he is rowing against the flow coming back his speed will be 2.5 + 1.5 = 4 kmph

OpenStudy (cwrw238):

suppose he travels x km up the stream so speeed = distance / time 1 = x / t1 where t1 = time taken coming back 4 = x / t2 where t2 = time taken to come back so we have x = t1 and x = 4t2 so t1 = 4t2 and he travels for total time of 7 hours can you continue from here? you need to find the values of x

OpenStudy (cwrw238):

* value of x